Solve each of the following problems by setting up and solving an algebraic equation. Suppose that Bab is paid two times his normal hourly rate for each hour worked excess of 40 hours in a week. Last week he earned $504 for 48 hours of work. What is his hourly age?
Answer by checkley77(12844)

40x+8*2x=504
40x+16x=504
56x=504
x=504/56
x=$9 is the hourly rate.
Proof:
40*9+8*18=504
360+144=504
504=504